OpenStudy (anonymous):

How does the graph change f(x) = x^2 to f(x +c)??? Narrower, wider, etc.

OpenStudy (anonymous):

\[f(x)=x^2\\ f(x+c)=(x+c)^2\] the graph will look just like \(y=x^2\) but shifted RIGHT \(c\) units if \(c<0\) i.e. if \(c\) is negative, shifted LEFT \(c\) units is \(c>0\) i.e. if \(c\) is positive